if a rectangle has on side as 12 and another as x. and a different rectangle has 2-x as one side and 16 as another, find the value of x so that the rectangles have the same area.
Answer by oberobic(2304)

Area = Length * Width
A = L*W
.
Rectangle 1:
.
L = 12
W = x
A = 12x
.
Rectangle 2:
.
L = 16
W = 2-x
A = 16(2-x)
A = 32-16x
.
12x = 32-16x
.
-4x = 32
x = -8
.
So this is impossible.
You cannot have a negative width.
There is something wrong with your problem statement.
.
Done.